# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

### v2

@@ Abstract
Acknowledgements. This work was supported by Welch Foundation grant F-2112.
Supplementary Material for ’Non-universal surface magnetoelectric response in antiferromagnetic topological insulators”
Appendix S1 Landau level representation
@@
$\hat{H}=\sum_{{\mathbf{k}}_{\perp},ij,\mu\gamma}\Big{[}\Big{(}\,(-1)^{i}\hbar v_{D}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}+m_{i}\sigma_{z}\Big{)}_{\mu,\gamma}\delta_{ij}+\Delta_{ij}(1-\delta_{ij})\delta_{\mu,\gamma}\Big{]}\hat{c}_{{\mathbf{k}}_{\perp},i,\mu}^{\dagger}\hat{c}_{{\mathbf{k}}_{\perp},j,\gamma}.$ (S1)
Here ${\mathbf{k}}_{\perp}\equiv(k_{x},k_{y},0)$ for stacking axis $\hat{\bm{z}}$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ label states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$ and $\mu,\gamma\in\{\uparrow,\downarrow\}$ label up and down spin components of $\sigma_{z}$, $\Delta_{ij}$ are intra- and inter-layer hybridization parameters, $v_{{}_{D}}$ is the velocity of the isolated Dirac cones, and $m_{i}$ are mass terms that account for exchange coupling of the Dirac-like states at surface $i$ to static local magnetic moments (taken aligned with the stacking direction) that can be present in each layer. We set $m_{i}=\sum_{l}J_{il}M_{l}$ with dimensionless magnetization $M_{l}=0$ if layer $l$ is non-magnetic layers and ±$\pm 1$ if layer $l$ is magnetic. As mentioned in the main text, we employ the usual prescription to account for an external magnetic field ${\mathbf{B}}=-B\hat{{\mathbf{z}}}$ in a ${\mathbf{k}}\cdot\bm{p}$ model: $\hbar{\mathbf{k}}_{\perp}\rightarrow-i\hbar(\partial/\partial x,\,\partial/\partial y,0)$ and choose the Landau gauge ${\mathbf{A}}=(0,-Bx,0)$. Under this recipe the Dirac Hamiltonian for each surface transforms as $H_{D}({\mathbf{k}}_{\perp})=\hbar v_{D}(\sigma_{x}k_{y}-\sigma_{y}k_{x})\rightarrow H_{D}(x,y)\equiv\hbar v_{D}\big{(}-\sigma_{x}(i\frac{\partial}{\partial y}+x/l_{B}^{2})+i\sigma_{y}\frac{\partial}{\partial x}\big{)}$ with $l_{B}\equiv\sqrt{\hbar c/eB}$. The total Hamiltonian that is obtained from employing this procedure in Eq. (S1) has translational symmetry in $\hat{{\mathbf{y}}}$, thus energy eigenfunctions can be written as
$\Psi_{E,q_{y}}(x,y)=\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\Phi_{E,q_{y}}(x)\equiv\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\Big{(}\phi_{E,q_{y},(i,\uparrow)}(x),\,\phi_{E,q_{y},(i,\downarrow)}(x)\Big{)}.$ (S2)
Since the only position dependence of the total Hamiltonian appears via the terms resembling $H_{D}$ for each surface, we focus on those contributions first. For an isolated surface of a non-magnetic layer the energy eigenfunctions satisfy $H_{D}^{2}(x,y)\Psi_{E,q_{y}}(x,y)=E^{2}\Psi_{E,q_{y}}(x,y)$ and $\Phi_{E,q_{y}}(x)=\big{(}\phi_{E,q_{y},\uparrow}(x),\,\phi_{E,q_{y},\downarrow}(x)\big{)}$, where
@@
$(\hbar v_{D})^{2}(-\partial_{x}+q_{y}-x/l_{B}^{2})(\partial_{x}+q_{y}-x/l_{B}^{2})\phi_{E,q_{y},\downarrow}(x)$ $=E^{2}\phi_{E,q_{y},\downarrow}(x).$ (S3)
With the substitutions $\tilde{x}\equiv l_{B}q_{y}-x/l_{B}$ and $\partial_{\tilde{x}}=-l_{B}\partial_{x}$, define the usual harmonic oscillator creation and annihilation differential operators $a^{\dagger}\equiv(\tilde{x}-\partial_{\tilde{x}})/\sqrt{2}$ and $a\equiv(\tilde{x}+\partial_{\tilde{x}})/\sqrt{2}$, which satisfy $[a,a^{\dagger}]=1$. Then, Eq. (S3) is recast as
$2\left(\frac{\hbar v_{D}}{l_{B}}\right)^{2}\begin{pmatrix}a^{\dagger}a&0\\ 0&aa^{\dagger}\end{pmatrix}\Phi_{E,q_{y}}(x)\equiv H^{2}_{D}(x,\partial_{x};q_{y})\Phi_{E,q_{y}}(x)=E^{2}\Phi_{E,q_{y}}(x).$ (S4)
@@
where $H_{n}(x)$ is the Hermite polynomial $H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$, and have corresponding eigenvalues $n$. These functions satisfy $a^{\dagger}\chi_{n}(\tilde{x})\propto\chi_{n+1}(\tilde{x})$ and $a\chi_{n+1}(\tilde{x})\propto\chi_{n}(\tilde{x})$ for integers $n\geq 0$. Using this we find that solutions of the eigenfunction-eigenvalue equation (S4) can be written in the general form
$\Phi_{E,q_{y}}(x)\equiv\Phi_{n}(\tilde{x})=\begin{cases}\big{(}C^{n}_{\uparrow}\chi_{n}(\tilde{x}),\,C^{n}_{\downarrow}\chi_{n-1}(\tilde{x})\big{)},&\text{ for }n\neq 0\\ \big{(}C^{0}_{\uparrow}\chi_{0}(\tilde{x}),\,0\big{)},&\text{ for }n=0\end{cases}$ (S6)
@@
for $C^{n}_{\mu}\in\mathbb{C}$ and the corresponding eigenvalues are $E_{n}^{2}=2n(\hbar v_{D}/l_{B})^{2}$. This implies $E_{n}=\pm\hbar\omega_{c}\sqrt{n}$ where $\omega_{c}\equiv\sqrt{2}v_{D}/l_{B}$. Indeed, these general forms can diagonalize the isolated non-magnetic surface Dirac Hamiltonian
@@
$H_{D}({\mathbf{k}}_{\perp})\rightarrow H_{D}(x,\partial_{x};q_{y})\equiv\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix}.$ (S7)
@@
In the total Hamiltonian (for a $N$ layer film) there is a contribution related to $H_{D}$ for each surface $i$ that couples $(\chi_{n}(\tilde{x}),0)$ only to $(0,\chi_{n-1}(\tilde{x}))$. A reasonable ansatz for an energy eigenfunction could therefore involve $\Phi_{E}(\tilde{x})$ of the form $\prod_{i=0}^{2N-1}\big{(}C^{n_{i}}_{(i,\uparrow)}\chi_{n_{i}}(\tilde{x}),\,C^{n_{i}}_{(i,\downarrow)}\chi_{n_{i}-1}(\tilde{x})\big{)}$, where $n_{i}\geq 0$ for all surfaces $i\in\{1,\ldots,2N-1\}$. However, since there are no other terms in (S1) that involve ${\mathbf{k}}_{\perp}$, there are no other terms in the Hamiltonian (after $\hbar{\mathbf{k}}_{\perp}\rightarrow-i\hbar(\partial/\partial x,\,\partial/\partial y)$) that involve $a$ or $a^{\dagger}$. That is, when acting on the ansatz eigenfunction the other terms in the Hamiltonian can map a component at surface $i$ into a component at surface $j$, but only in such a way that the new component at surface $j$ is in the space spanned by $\chi_{n_{i}}(\tilde{x})$. In our model the states at each surface are coupled to those at two surfaces adjacent to it (via $\Delta_{\text{S}}$ and $\Delta_{\text{D}}$), thus the ansatz is valid only if $n_{1}=\ldots=n_{2N-1}\equiv n$ for $n\geq 0$. (We sometimes adopt the notation $\mathcal{N}$ for this $n$.) The total energy eigenfunctions can therefore be written in the general form
@@
$\Psi_{n,q_{y}}(x,y)=\begin{cases}\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)}\chi_{n}(\tilde{x}),\,C^{n}_{(i,\downarrow)}\chi_{n-1}(\tilde{x})\big{)},&\text{ for }n\neq 0\\ \frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\big{(}C^{0}_{(i,\uparrow)}\chi_{0}(\tilde{x}),\,0\big{)},&\text{ for }n=0\end{cases}.$ (S8)
S1.1 Multi-layer antiferromagnetic materials
@@
In this section we specify the action of the Hamiltonian (for a material consisting of $N$ layers) in the subspace defined by LL index $n$ via a matrix acting on a collection of eigenfunction components $C^{n}_{(i,\mu)}$ introduced in Eq. (S8). For $n\neq 0$ the Hamiltonian matrix acts on the collection of components $\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)},\,C^{n}_{(i,\downarrow)}\big{)}$. The bulk ($N\rightarrow\infty$) multi-layer antiferromagnetic Hamiltonian matrix for $n\neq 0$ is
@@
$H_{\text{bulk}}^{n}=\left(\begin{array}[]{cccccccc}J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&\Delta_{\text{D}}e^{-2ik_{z}d}&0\\ \sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{S}}&0&0&0&\Delta_{\text{D}}e^{-2ik_{z}d}\\ \Delta_{\text{S}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{S}}\\ \Delta_{\text{D}}e^{2ik_{z}d}&0&0&0&\Delta_{\text{S}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}\\ 0&\Delta_{\text{D}}e^{2ik_{z}d}&0&&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}\\ \end{array}\right).$ (S9)
@@
$H_{\text{2L}}^{n}=\left(\begin{array}[]{cccccccc}J_{\text{S}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&0&0\\ \sqrt{n}\hbar\omega_{c}&-J_{\text{S}}&0&\Delta_{\text{S}}&0&0&0&0\\ \Delta_{\text{S}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{S}}\\ 0&0&0&0&\Delta_{\text{S}}&0&-J_{\text{S}}&\sqrt{n}\hbar\omega_{c}\\ 0&0&0&&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}\\ \end{array}\right).$ (S10)
@@
$MH_{\text{2L}}^{n}M=\left(\begin{array}[]{cccccccc}-J_{\text{S}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&0&0\\ \sqrt{n}\hbar\omega_{c}&J_{\text{S}}&0&\Delta_{\text{S}}&0&0&0&0\\ \Delta_{\text{S}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{S}}\\ 0&0&0&0&\Delta_{\text{S}}&0&J_{\text{S}}&\sqrt{n}\hbar\omega_{c}\\ 0&0&0&&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}\\ \end{array}\right).$ (S11)
@@
$H^{n=0}_{\text{bulk}}=\left(\begin{array}[]{cccc}J_{\text{S}}-J_{\text{D}}&\Delta_{\text{S}}&0&\Delta_{\text{D}}e^{-2ik_{z}d}\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{D}}&0\\ 0&\Delta_{\text{D}}&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ \Delta_{\text{D}}e^{2ik_{z}d}&0&\Delta_{\text{S}}&-J_{\text{S}}+J_{\text{D}}\\ \end{array}\right).$ (S12)
In the case of a 2-layer thin film the Hamiltonian matrix is
$H^{n=0}_{\text{2L}}=\left(\begin{array}[]{cccc}J_{\text{S}}&\Delta_{\text{S}}&0&0\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{D}}&0\\ 0&\Delta_{\text{D}}&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (S13)
@@
At $k_{z}=0$ the energy eigenvalues of $H^{n=0}_{\text{bulk}}$ (Eq. (S12)) are:
@@
$E_{2}$ $=-E_{3}=\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}-\Delta_{\text{S}},$ (S14)
with corresponding (non-normalized) eigenvectors:
@@
$\Psi_{4}$ $=\left(\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-1,1\right).$ (S15)
@@
S1.2 Bulk polarization of the electronic ground state in the low-energy model

### v4

@@ VI Acknowledgements
Work at the University of Texas at Austin was supported by the Robert A. Welch Foundation under Grant Welch F-2112.
Appendix A Landau level representation
@@
$\hat{H}=\sum_{{\mathbf{k}}_{\perp},ij,\mu\gamma}\Big{[}\Big{(}\,(-1)^{i}\hbar v_{D}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}+m_{i}\sigma_{z}\Big{)}_{\mu,\gamma}\delta_{ij}+\Delta_{ij}(1-\delta_{ij})\delta_{\mu,\gamma}\Big{]}\hat{c}_{{\mathbf{k}}_{\perp},i,\mu}^{\dagger}\hat{c}_{{\mathbf{k}}_{\perp},j,\gamma}.$ (10)
Here ${\mathbf{k}}_{\perp}\equiv(k_{x},k_{y},0)$ for stacking axis $\hat{\bm{z}}$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ label states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$ and $\mu,\gamma\in\{\uparrow,\downarrow\}$ label up and down spin components of $\sigma_{z}$, $\Delta_{ij}$ are intra- and interlayer hybridization parameters, $v_{{}_{D}}$ is the velocity of the isolated Dirac cones, and $m_{i}$ are mass terms that account for exchange coupling of the Dirac-like states at surface $i$ to static local magnetic moments (taken aligned with the stacking direction) that can be present in each layer. We set $m_{i}=\sum_{l}J_{il}M_{l}$ with dimensionless magnetization $M_{l}=0$ if layer $l$ is nonmagnetic layers and ±$\pm 1$ if layer $l$ is magnetic. As mentioned in the main text, we employ the usual prescription to account for an external magnetic field ${\mathbf{B}}=-B\hat{{\mathbf{z}}}$ in a ${\mathbf{k}}\cdot\bm{p}$ model: $\hbar{\mathbf{k}}_{\perp}\rightarrow-i\hbar(\partial/\partial x,\,\partial/\partial y,0)$ and choose the Landau gauge ${\mathbf{A}}=(0,-Bx,0)$. Under this recipe the Dirac Hamiltonian for each surface transforms as $H_{D}({\mathbf{k}}_{\perp})=\hbar v_{D}(\sigma_{x}k_{y}-\sigma_{y}k_{x})\rightarrow H_{D}(x,y)\equiv\hbar v_{D}\big{(}-\sigma_{x}(i\frac{\partial}{\partial y}+x/l_{B}^{2})+i\sigma_{y}\frac{\partial}{\partial x}\big{)}$ with $l_{B}\equiv\sqrt{\hbar c/eB}$. The total Hamiltonian that is obtained from employing this procedure in Eq. (10) has translational symmetry in $\hat{{\mathbf{y}}}$, thus energy eigenfunctions can be written as
$\Psi_{E,q_{y}}(x,y)=\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\Phi_{E,q_{y}}(x)\equiv\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\Big{(}\phi_{E,q_{y},(i,\uparrow)}(x),\,\phi_{E,q_{y},(i,\downarrow)}(x)\Big{)}.$ (11)
Since the only position dependence of the total Hamiltonian appears via the terms resembling $H_{D}$ for each surface, we focus on those contributions first. For an isolated surface of a nonmagnetic layer the energy eigenfunctions satisfy $H_{D}^{2}(x,y)\Psi_{E,q_{y}}(x,y)=E^{2}\Psi_{E,q_{y}}(x,y)$ and $\Phi_{E,q_{y}}(x)=\big{(}\phi_{E,q_{y},\uparrow}(x),\,\phi_{E,q_{y},\downarrow}(x)\big{)}$, where
@@
$(\hbar v_{D})^{2}(-\partial_{x}+q_{y}-x/l_{B}^{2})(\partial_{x}+q_{y}-x/l_{B}^{2})\phi_{E,q_{y},\downarrow}(x)$ $=E^{2}\phi_{E,q_{y},\downarrow}(x).$ (12)
With the substitutions $\tilde{x}\equiv l_{B}q_{y}-x/l_{B}$ and $\partial_{\tilde{x}}=-l_{B}\partial_{x}$, define the usual harmonic oscillator creation and annihilation differential operators $a^{\dagger}\equiv(\tilde{x}-\partial_{\tilde{x}})/\sqrt{2}$ and $a\equiv(\tilde{x}+\partial_{\tilde{x}})/\sqrt{2}$, which satisfy $[a,a^{\dagger}]=1$. Then, Eq. (12) is recast as
$2\left(\frac{\hbar v_{D}}{l_{B}}\right)^{2}\begin{pmatrix}a^{\dagger}a&0\\ 0&aa^{\dagger}\end{pmatrix}\Phi_{E,q_{y}}(x)\equiv H^{2}_{D}(x,\partial_{x};q_{y})\Phi_{E,q_{y}}(x)=E^{2}\Phi_{E,q_{y}}(x).$ (13)
@@
where $H_{n}(x)$ is the Hermite polynomial $H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$, and have corresponding eigenvalues $n$. These functions satisfy $a^{\dagger}\chi_{n}(\tilde{x})\propto\chi_{n+1}(\tilde{x})$ and $a\chi_{n+1}(\tilde{x})\propto\chi_{n}(\tilde{x})$ for integers $n\geq 0$. Using this we find that solutions of the eigenfunction-eigenvalue equation (13) can be written in the general form
$\Phi_{E,q_{y}}(x)\equiv\Phi_{n}(\tilde{x})=\begin{cases}\big{(}C^{n}_{\uparrow}\chi_{n}(\tilde{x}),\,C^{n}_{\downarrow}\chi_{n-1}(\tilde{x})\big{)},&\text{ for }n\neq 0\\ \big{(}C^{0}_{\uparrow}\chi_{0}(\tilde{x}),\,0\big{)},&\text{ for }n=0\end{cases}$ (15)
@@
for $C^{n}_{\mu}\in\mathbb{C}$ and the corresponding eigenvalues are $E_{n}^{2}=2n(\hbar v_{D}/l_{B})^{2}$. This implies $E_{n}=\pm\hbar\omega_{c}\sqrt{n}$ where $\omega_{c}\equiv\sqrt{2}v_{D}/l_{B}$. Indeed, these general forms can diagonalize the isolated nonmagnetic surface Dirac Hamiltonian
@@
$H_{D}({\mathbf{k}}_{\perp})\rightarrow H_{D}(x,\partial_{x};q_{y})\equiv\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix}.$ (16)
@@
In the total Hamiltonian (for a $N$ layer film) there is a contribution related to $H_{D}$ for each surface $i$ that couples $(\chi_{n}(\tilde{x}),0)$ only to $(0,\chi_{n-1}(\tilde{x}))$. A reasonable ansatz for an energy eigenfunction could therefore involve $\Phi_{E}(\tilde{x})$ of the form $\prod_{i=0}^{2N-1}\big{(}C^{n_{i}}_{(i,\uparrow)}\chi_{n_{i}}(\tilde{x}),\,C^{n_{i}}_{(i,\downarrow)}\chi_{n_{i}-1}(\tilde{x})\big{)}$, where $n_{i}\geq 0$ for all surfaces $i\in\{1,\ldots,2N-1\}$. However, since there are no other terms in Eq. (10) that involve ${\mathbf{k}}_{\perp}$, there are no other terms in the Hamiltonian (after $\hbar{\mathbf{k}}_{\perp}\rightarrow-i\hbar(\partial/\partial x,\,\partial/\partial y)$) that involve $a$ or $a^{\dagger}$. That is, when acting on the ansatz eigenfunction the other terms in the Hamiltonian can map a component at surface $i$ into a component at surface $j$, but only in such a way that the new component at surface $j$ is in the space spanned by $\chi_{n_{i}}(\tilde{x})$. In our model the states at each surface are coupled to those at two surfaces adjacent to it (via $\Delta_{\text{S}}$ and $\Delta_{\text{D}}$), thus the ansatz is valid only if $n_{1}=\ldots=n_{2N-1}\equiv n$ for $n\geq 0$. (We sometimes adopt the notation $\mathcal{N}$ for this $n$.) The total energy eigenfunctions can therefore be written in the general form
@@
$\Psi_{n,q_{y}}(x,y)=\begin{cases}\frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)}\chi_{n}(\tilde{x}),\,C^{n}_{(i,\downarrow)}\chi_{n-1}(\tilde{x})\big{)},&\text{ for }n\neq 0\\ \frac{e^{iq_{y}y}}{\sqrt{L_{y}}}\prod_{i=0}^{2N-1}\big{(}C^{0}_{(i,\uparrow)}\chi_{0}(\tilde{x}),\,0\big{)},&\text{ for }n=0\end{cases}.$ (17)
A.1 Multi-layer antiferromagnetic materials
@@
In this section we specify the action of the Hamiltonian (for a material consisting of $N$ layers) in the subspace defined by LL index $n$ via a matrix acting on a collection of eigenfunction components $C^{n}_{(i,\mu)}$ introduced in Eq. (17). For $n\neq 0$ the Hamiltonian matrix acts on the collection of components $\prod_{i=0}^{2N-1}\big{(}C^{n}_{(i,\uparrow)},\,C^{n}_{(i,\downarrow)}\big{)}$. The bulk ($N\rightarrow\infty$) multi-layer antiferromagnetic Hamiltonian matrix for $n\neq 0$ is
@@
$H_{\text{bulk}}^{n}=\left(\begin{array}[]{cccccccc}J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&\Delta_{\text{D}}e^{-2ik_{z}d}&0\\ \sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{S}}&0&0&0&\Delta_{\text{D}}e^{-2ik_{z}d}\\ \Delta_{\text{S}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{S}}\\ \Delta_{\text{D}}e^{2ik_{z}d}&0&0&0&\Delta_{\text{S}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}\\ 0&\Delta_{\text{D}}e^{2ik_{z}d}&0&0&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}\\ \end{array}\right).$ (18)
@@
$H_{\text{2L}}^{n}=\left(\begin{array}[]{cccccccc}J_{\text{S}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&0&0\\ \sqrt{n}\hbar\omega_{c}&-J_{\text{S}}&0&\Delta_{\text{S}}&0&0&0&0\\ \Delta_{\text{S}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{S}}\\ 0&0&0&0&\Delta_{\text{S}}&0&-J_{\text{S}}&\sqrt{n}\hbar\omega_{c}\\ 0&0&0&0&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}\\ \end{array}\right).$ (19)
@@
$MH_{\text{2L}}^{n}M=\left(\begin{array}[]{cccccccc}-J_{\text{S}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0&0&0&0&0\\ \sqrt{n}\hbar\omega_{c}&J_{\text{S}}&0&\Delta_{\text{S}}&0&0&0&0\\ \Delta_{\text{S}}&0&-J_{\text{S}}+J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&J_{\text{S}}-J_{\text{D}}&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&J_{\text{S}}-J_{\text{D}}&\sqrt{n}\hbar\omega_{c}&\Delta_{\text{S}}&0\\ 0&0&0&\Delta_{\text{D}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}+J_{\text{D}}&0&\Delta_{\text{S}}\\ 0&0&0&0&\Delta_{\text{S}}&0&J_{\text{S}}&\sqrt{n}\hbar\omega_{c}\\ 0&0&0&0&0&\Delta_{\text{S}}&\sqrt{n}\hbar\omega_{c}&-J_{\text{S}}\\ \end{array}\right).$ (20)
@@
$H^{n=0}_{\text{bulk}}=\left(\begin{array}[]{cccc}J_{\text{S}}-J_{\text{D}}&\Delta_{\text{S}}&0&\Delta_{\text{D}}e^{-2ik_{z}d}\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{D}}&0\\ 0&\Delta_{\text{D}}&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ \Delta_{\text{D}}e^{2ik_{z}d}&0&\Delta_{\text{S}}&-J_{\text{S}}+J_{\text{D}}\\ \end{array}\right).$ (21)
In the case of a two-layer thin film the Hamiltonian matrix is
$H^{n=0}_{\text{2L}}=\left(\begin{array}[]{cccc}J_{\text{S}}&\Delta_{\text{S}}&0&0\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{D}}&0\\ 0&\Delta_{\text{D}}&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (22)
@@
At $k_{z}=0$ the energy eigenvalues of $H^{n=0}_{\text{bulk}}$ (Eq. (21)) are:
@@
$E_{2}$ $=-E_{3}=\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}-\Delta_{\text{S}},$ (23)
with corresponding (nonnormalized) eigenvectors:
@@
$\Psi_{4}$ $=\left(\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-\frac{J_{\text{S}}-J_{\text{D}}-\sqrt{\Delta_{\text{D}}^{2}+(J_{\text{S}}-J_{\text{D}})^{2}}}{\Delta_{\text{D}}},-1,1\right).$ (24)
@@
A.2 Bulk polarization of the electronic ground state in the low-energy model